question......write an equation in point slope form and slope intercept form of the line passing through (-12,1) and perpendicular to the line whose equation is 6x-y-4=0
Answer by stanbon(75887)

question......write an equation in point slope form and slope intercept form of the line passing through (-12,1) and perpendicular to the line whose equation is 6x-y-4=0
-----------------------------
The given equation has a slope:
y=6x-4, so the slope is 6
-------------
The perpendicular line must have slope of -1/6.
----------
You want the equation of the line with slope=-1/6 passing though (-12,1)
Use the point-slope form to get:
y-1 = (-1/6)(x+12)
y=(-1/6)x-1
That is the equation you want.
Cheers,
Stan H.